What is the volume of Oxygen required to produce 300cm3 of SO3 at room temperature and pressure
1
Expert's answer
2020-04-06T11:45:11-0400

2S+3O2=2SO3


300 ml = 0,3l


n(SO3)=V(SO3)/Vm=0,3/22,4=0.013mol

n(O2)=0.013*3/2=0.0195mol


V(O2)=0.0195×22,4=0.437l=437ml


Answer 437ml